what is devidendo and componendo?]
Pls explain me is it is used as shortcut method or it is the only method.
 

Grade:11

1 Answers

Vikas TU
6 years ago
Componendo govern basically states 
In the event that a/b=c/d, at that point a+b/b=c+d/d 
Dividendo govern states: 
On the off chance that a/b=c/d, at that point a−b/b=c−d/d 
Componendo and Dividendo run states 
In the event that a/b=c/d, at that point a+b/a−b=c+d/c−d .